# Discovering Knowledge-Critical Subnetworks in Pretrained Language Models

### v1

@@ Appendix D Additional Subnetwork Discovery Results
Appendix E Additional Parameter Removal and Addition Analysis
Figure 3: Removing and adding parameters to the remaining model, averaged over five seeds, with standard deviation depicted as the filled area around the average curves. The $x$-axis is the removed subnetwork sparsity. The $y$-axis is the $\Delta$ PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$) for the different datasets. Vertical dashed lines show the original sparsity of the critical subnetwork. The darker curve is the outcome starting from the critical subnetwork, whereas the lighter curve is from a randomly masked model at the same sparsity.
We include additional KG results for removing and adding parameters to the remaining model (Figure 3), specifically for location and communication. As mentioned in §6.3, the TargetKG recovery extent seems to depend on the degree the knowledge was suppressed in the first place. If the model forgot to express the knowledge by a significant amount, removing further parameters decreases TargetKG $\Delta$ PPL slowly and not as significantly when the critical subnetwork is compared to the randomly masked baseline (e.g., representation and location). On the other hand, the communication-critical subnetwork was at sparsity $\approx$99with a TargetKG perplexity difference of 166, and further removing parameters decreases the perplexity difference on average to 60 at 92% sparsity, indicating that its suppression may have overfit to a particular configuration. However, we note that even once it’s $\Delta$ PPL drops to 60, this difference is still higher than a randomly masked baseline average.
Appendix F Additional Details on Downstream Task Transfer
To learn a mask for a set of ConceptNet relations, we need to verbalize them with a relation-specific prompt. As described in §6.4, CommonsenseQA questions are not explicitly annotated with a relation. However, they were constructed with ConceptNet such that each question’s head concept relates to four of the tail answers with the same relation. This does not apply to the fifth answer, as crowd workers created them. Therefore, to retrieve the relations, we iterate through the questions and check if any relations with the question head concept and correct tail answer exist in the LAMA and Commonsense Knowledge Base Completion subsets of ConceptNet (Li et al., 2016; Petroni et al., 2019). If it does and has only one relation, we choose that relation. If it has multiple relations, we take the union of relations between the head concept and the distractor tail answers, and intersect that set with the correct tail relations. If the intersection is a set larger than one element, we choose one relation at random. Out of the 1221 test questions, only 572 have a single-token correct answer, and we could only find the corresponding relation to 363 questions, which is our filtered test set.
For the MCQA head, we use the Huggingface Double Heads model. [^6] https://huggingface.co/docs/transformers In addition to the language modeling head, this model adds a parallel multiple-choice classification head. The MCQA head takes as input the last sequence output. To finetune the MCQA model, we use three kinds of fine-tuning. The first one is Head Tuning, in which the model parameters are frozen, but the MCQA head is not. The second method is LoRA (Hu et al., 2022), which is a parameter-efficient finetuning method. Similar to the head tuning method, LoRA freezes the model parameters and instead inserts trainable rank decomposition parameters in each transformer layer. We use a rank of 16 for all LoRA experiments. Finally, we also try Full Finetuning, in which all model parameters are tuned. To remove a subnetwork, we manually set the knowledge-critical parameters to 0. Therefore, the value of these parameters can change during full finetuning.
Table 15: Accuracy on downstream CommonsenseQA task, averaged over three seeds. Ours refers to removing the critical subnetwork. Random refers to removing a random subnetwork at the same sparsity as the critical subnetwork.
@@ Appendix F Additional Details on Downstream Task Transfer
In addition, we also verify whether learning a mask for one randomly selected half of the filtered test set (Half 1) corrupts downstream task transfer for the half (Half 2), where there are no triplet overlaps. We find in Table 15 that, on average, the accuracy on the triplets the mask was trained for is less by 3.6% than the held-out half.
Appendix G Paraphrase Generalization
To make sure that the remaining model is not overfit to the specific prompt we used during training and that TargetKG knowledge is removed with other prompting styles, we evaluate the remaining models on paraphrases of the knowledge triplet. We run an experiment where we evaluate $\Delta$ PPL of TargetKG and ControlKG for all prompt styles, excluding the one used in the training process, which had the lowest tail-token perplexity on the pretrained model. This results in 20 other distinct templates that differ either by the relation token, by the entity articles, or with a space put in front of the sentence. Note that the average perplexity of the original model on the worse paraphrases is 3231 for TargetKG and 2368 for ControlKG averaged across three KGs (representation, location, and communication).
Table 16: Success metrics on paraphrases.
Knowledge TargetKG ControlKG
Graph $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$)
communication 492.4 3.0
location 684.9 -33.0
representation 916.4 -6.9
The success metrics do not change when using other prompt styles, as seen in Table 16 compared to Table 1. We note that the model perplexity on ControlKG paraphrases is a bit lower than the format used during training, likely because the default perplexity is higher on other templates. Because it is higher than the original templates for ControlKG, the model may generalize its maintenance of ControlKG to a greater degree on these suboptimal templates. In summary, our methods generalizes to also suppress alternate prompts of TargetKG.
Appendix H Seed-Based Analysis
@@
We investigate the stability of subnetwork discovery under seed-based variance. We also explore whether composing subnetworks from different seeds could increase the suppression effect while still fulfilling the rest of the success criteria.
@@
Prior work shows that subnetworks identified under distinct random seeds may differ with a large variance Csordás et al. (2021). We inspect how subnetworks from three best checkpoints for three random seeds overlap for an individual TargetKG. We use Jaccard similarity, or intersection over union (IoU), as the overlap metric. In Figure 4, we plot a Venn diagram of parameter overlap for each knowledge graph. We can see that, on average, when using IoU, only around 3.7% of the unioned subnetwork parameters overlap across the three seeds (3.76% for location, 3.8% for communication, and 3.5% for representation), meaning the subnetworks identified under different random seeds vary and that models may encode multiple knowledge-critical subnetworks for the same concept. Across layers, the IoU is also low with a slightly higher overlap for the final attention layer masks ($\approx$10%) as shown in Figure 5.
Figure 4: Venn diagrams for parameter overlap of three subnetworks identified under three different random seeds, for each KG representation, location, and communication.
Figure 5: Jaccard similarity for the different seeds of KGs representation, location, and communication. The brighter the color, the higher the Intersection over Union.
@@
Table 17: Composing subnetworks across seeds, averaged across three KGs. Original stands for the individual subnetwork removal average across the same three seeds and KGs.
@@
Pattern (normal-$\uparrow$) $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
@@
Union 96.9 4013.3 31.0 5.3
Floral 99.5 84.8 8.9 1.4
Intersect 99.9 10.6 6.6 2.4
In Table 17, when we combine the masks and remove them, we see that the maintenance datasets ControlKG and ControlLM are least affected by using the floral and intersection composition, perhaps due to their low sparsity. Removing the union of the subnetworks increases the TargetKG perplexity difference larger than the original results. However, in all three cases, combining the subnetworks and removing them has more side effects on maintenance datasets than using an individual seed’s subnetwork, as seen in the original results. A future idea could be to continue optimizing for the subnetwork mask by initializing it as the union of the subnetworks to see if more robust removal can be achieved.

### v2

@@ Appendix D Additional Subnetwork Discovery Results
communication.n.02 99.3 [99.2, 99.5] 1.0 [0.3, 1.7] -0.3 [-0.3, -0.2] 0.1 [0.1, 0.1]
location.n.01 99.4 [99.1, 99.7] 1.3 [-0.7, 3.3] -0.6 [-1.2, 0.0] 0.1 [0.1, 0.1]
XL representation.n.02 99.2 [99.1, 99.2] 2.4 [-0.5, 5.4] 1.2 [0.6, 1.9] 0.2 [0.2, 0.2]
Average 99.3 [99.2, 99.4] 1.6 [1.0, 2.4] 0.1 [-0.6, 1.2] 0.1 [0.1, 0.2]
Table 18: Subnetwork discovery results on larger randomly masked models per KG with weight masking, averaged over two seeds.
Minimum & Maximum Boundaries
In addition to the average $\Delta$PPL and $\Delta$Rank presented in Table 2, we add minimum and maximum boundaries to all of the results in Table 13 and 15. We also provide log probability differences $\Delta$LogProb similar to how $\Delta$PPL is calculated. We observe in Table 15 the same trend as $\Delta$PPL. On average, removing the subnetwork increases the rank of the gold tail token and decreases the log probability. In contrast, the randomly masked baseline does not increase the TargetKG rank significantly and does not maintain ControlKG rank to the same extent as the critical subnetwork.
Model Scale
Randomly Masked Baseline
Appendix E Additional Details on Downstream Task Transfer
@@
Table 19: Accuracy on downstream CommonsenseQA task with GPT2-small and weight masking, averaged over three seeds. Ours refers to removing the critical subnetwork. Random refers to removing a random subnetwork at the same sparsity as the critical subnetwork.
To learn a mask for a set of ConceptNet relations, we need to verbalize them with a relation-specific prompt. As described in §6.2, CommonsenseQA questions are not explicitly annotated with a relation. However, they were constructed with ConceptNet such that each question’s head concept relates to four of the tail answers with the same relation. This does not apply to the fifth answer, as crowd workers created them. Therefore, to retrieve the relations, we iterate through the questions and check if any relations with the question head concept and correct tail answer exist in the LAMA and Commonsense Knowledge Base Completion subsets of ConceptNet Li et al. (2016); Petroni et al. (2019). If it does and has only one relation, we choose that relation. If it has multiple relations, we take the union of relations between the head concept and the distractor tail answers and intersect that with the correct tail triplets. If the intersection is a set larger than one element, we choose one relation at random. Out of the 1221 test questions, only 572 have a single-token correct answer, and we could only find the corresponding relation to 363 questions, which is our filtered test set.
For the MCQA head, we use the Huggingface Double Heads model. [^15] https://huggingface.co/docs/transformers In addition to the language modeling head, this model adds a parallel multiple-choice classification head. The MCQA head takes as input the last sequence output. To finetune the MCQA model, we use three kinds of fine-tuning. The first one is Head Tuning, in which the model parameters are frozen, but the MCQA head is not. The second method is LoRA Hu et al. (2022), which is a parameter-efficient finetuning method. Similar to the head tuning method, LoRA freezes the model parameters and instead inserts trainable rank decomposition parameters in each transformer layer. We use a rank of 16 for all LoRA experiments. Finally, we also try Full Finetuning, in which all model parameters are tuned. To remove a subnetwork, we manually set the knowledge-critical parameters to 0. Therefore, the value of these parameters can change during full finetuning.
@@
In addition, we also verify whether learning a mask for one randomly selected half of the filtered test set (Half 1) corrupts downstream task transfer for a distinct half (Half 2), where there are no triplet overlaps. We find in Table 19 that, on average, the accuracy on the triplets the mask was trained for is less by 3.6% than the held-out half.
Appendix F Alternative Objective: Is expressing knowledge enough to be a knowledge-critical subnetwork?
Objective Sparsity TargetKG ControlKG ControlLM
Combination ($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
Table 20: Expression loss study with GPT2-small and weight masking, averaged across three KGs and two seeds. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
Appendix G Spurious Subnetworks Test
Figure 2: Removing and adding parameters to the remaining GPT2-small model, averaged over five seeds, with standard deviation depicted as the filled area around the average curves. The $x$-axis is the removed subnetwork sparsity. The $y$-axis is the $\Delta$PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$) for the different datasets. Vertical dashed lines show the original sparsity of the critical subnetwork. The darker curve is the outcome starting from the critical subnetwork, whereas the lighter curve is from a randomly masked model at the same sparsity.
We hypothesize that a spurious subnetwork would cause the remaining network from which it was removed to re-gain the ability to express TargetKG if the subnetwork was randomly expanded (i.e., $\Delta$PPL on TargetKG would drop as more parameters are removed from $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$). Meanwhile, if removing the critical subnetwork is not a spurious solution to suppress the TargetKG, then the remaining model would generally still fail to recognize TargetKG, even as more parameters were randomly removed, leading $\Delta$PPL to rise or stay the same. To verify this hypothesis, we remove further parameters from the remaining model. Starting from the knowledge-critical subnetwork sparsity, we randomly remove parameters at intervals of 0.5%. We run this iterative process of removing parameters with five different random seeds. We also test whether the mask has found a spurious solution to achieve the maintenance criteria by adding back parameters, though with smaller intervals of 0.1%, as the starting sparsity level is typically high.
In Figure 2, we observe that removing more parameters in small amounts does not significantly recover expressing TargetKG. As a baseline, we plot the effect on $\Delta$PPL of removing further parameters from remaining models with randomly removed subnetworks of the same sparsity. Interestingly, for the maintenance datasets, $\Delta$PPL for both datasets increases as we remove parameters from the remaining model. When we add back parameters, we do not see a linear recovery to $\Delta$PPL $=0$. Instead, we observe an initial phase of increase followed by a phase of decrease as the model returns to its original state (i.e., a $\Delta$PPL of zero at 100% sparsity). This effect can be explained by the fact that our subnetwork had been optimized to keep these abilities, and has been slightly overfit for maintenance, though not for suppression. Thus, randomly adding parameters back yields new sub-optimal pathways that corrupt the model’s original distribution.
Appendix H Structural Analysis
In this section, we investigate the structure of the removed knowledge-critical subnetworks by looking at their relative density across different layer types (Figure 3), and more specifically, across different attention heads (Figure 5) and the $W_{q}$, $W_{k}$, and $W_{v}$ matrices in attention sublayers (Figure 6. The density is calculated relatively, meaning according to the particular sublayer’s size. The model used is GPT2-small.
Layer depth-wise, we observe that the subnetwork is consistently most dense around the first and final masked transformer blocks, which are layers 7 and 12 in Figure 3. Specifically, layer type-wise, we find that knowledge-critical subnetworks are most dense in the attention sublayers for layer 7 and layer 12 (Attn-Out and Attn-$W_{q},W_{k},W_{v}$).
In addition, we have not found any complete columns or rows that were dense in the critical subnetworks. This means no input or output neuron features get completely removed when the critical subnetwork is removed. Therefore, the masked region may not be working to zero-out the knowledge by turning specific features off, which would counter the prevailing view that neuron-level changes are necessary for mechanistic interventions Dai et al. (2022); Meng et al. (2022).
When we investigated attention heads and $W_{q}$, $W_{k}$, and $W_{v}$ masks in detail for 3 KGs and 3 seeds, we found that head 10 in layer 7, and heads 1 and 9 in layer 12 are significantly dense. Moreover, the $W_{v}$ mask is consistently the most dense across the three attention $W_{q}$, $W_{k}$, and $W_{v}$ masks. Therefore, while the subnetworks do not have a significant IoU, as demonstrated by the seed-based (Appendix I) and the KG-based analyses (Appendix J), the subnetworks still tend to be dense in similar layer types at similar layer depths.
Appendix I Random Seed-Based Analysis
@@
We investigate the stability of subnetwork discovery under random seed variance for GPT2-small. We also explore whether composing subnetworks from different seeds could increase the suppression effect while still fulfilling the rest of the success criteria.
@@
Prior work shows that subnetworks identified under distinct random seeds may differ with a large variance Csordás et al. (2021). We inspect how subnetworks from the best checkpoints for three random seeds overlap for an individual TargetKG. We use Jaccard similarity, or intersection over union (IoU), as the overlap metric. In Figure 8, we plot a Venn diagram of parameter overlap for each knowledge graph. We can see that, on average, when using IoU, only around 3.7% of the unioned subnetwork parameters overlap across the three seeds (3.76% for location, 3.8% for communication, and 3.5% for representation), meaning the subnetworks identified under different random seeds vary, which complies with prior works’ analysis. Across layers, the IoU is also similarly low with a higher overlap for the final attention layer masks ($\approx$10%) as shown in Figure 10.
@@
In Table 5, we observe that removing the intersection and floral intersection of the subnetworks does not increase TargetKG $\Delta$PPL. On the other hand, removing the union of the subnetworks increases the TargetKG perplexity difference significantly larger than the original results. However, combining the subnetworks and removing them increases $\Delta$PPL on maintenance datasets more than using an individual seed’s subnetwork, as seen in the original results. We note that the increase in the $\Delta$PPL on maintenance datasets matches the increase we get when removing an equally sparse random subnetwork (see Table 2). Therefore, it may be possible to naively combine subnetworks; however, they may not guarantee the maintenance criteria to the same extent. A future idea could be to continue optimizing for the subnetwork mask by initializing it as the union of the subnetworks to see if more robust suppression can be achieved.
Appendix J Knowledge-Based Analysis
This section examines the overlap of subnetworks across different KGs for the same seed with GPT2-small. This contrasts with the previous section that studies the overlap of subnetworks across different seeds for the same KG. Similarly, we use Jaccard similarity, or intersection over union (IoU), as the overlap metric. We also explore whether composing subnetworks for different KGs from the same seed could suppress all of the TargetKGs.
Knowledge-based Variance
In Figure 12, we plot a Venn diagram of parameter overlap for each seed across different TargetKGs. On average, when using IoU, only around 3.56% of the unioned subnetwork parameters overlap across the three seeds (4.08% for seed 735, 4.01% for seed 1318, and 2.65% for seed 84). Across layers, the IoU is also similarly low with a significantly higher overlap for the final attention layer masks ($\approx$12%) as shown in Figure 13.
Subnetwork Composition
We combine masks of three KGs for the same seed in their intersection, their floral intersection (intersection unioned with each intersection of two KGs), and overall union to measure the effect on $\Delta$PPL for TargetKG, ControlKG, and ControlLM. We average the results over three seeds (735, 1318, and 84).
Similar to the findings in composing subnetworks for different seeds, Table 21 shows that composing subnetworks for different KGs increases the $\Delta$PPL on TargetKG when using their union. However, removing the union of the subnetworks also has higher perplexity differences on maintenance datasets than using an individual KG’s subnetwork, as seen in the original results. Once again, this $\Delta$PPL increase on the maintenance datasets matches the difference we would observe using an equally sparse random subnetwork. Therefore, while subnetworks of different KGs may be composable to fortify the suppression effect, they may not guarantee the maintenance criteria to the same extent as the individual subnetworks.
@@
Pattern ($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
@@
Union 96.9 1984.4 44.9 4.7
Floral 99.5 4.9 4.5 1.1
Intersection 99.9 7.9 3.7 2.1
Table 21: Composing subnetworks across KGs with GPT2-small and weight masking, averaged across three seeds. Original stands for the individual subnetwork removal average across the same three seeds and KGs.
Figure 3: Average module mask density with weight masking, for different KGs ( representation, location, and communication) and seeds. Reported in percentage (%). The brighter the color, the higher the removed mask density.
Figure 4: Average module mask density with neuron masking, for different KGs ( representation, location, and communication) and seeds. Reported in percentage (%). The brighter the color, the higher the removed mask density.
Figure 5: Density percentage (%) of different heads across different attention layers for weight masking. Each row represents a different KG and each column is a different seed.
Figure 6: Density percentage (%) of $W_{q}$, $W_{k}$, and $W_{v}$ masks in attention layers for weight masking. Each row represents a different KG and each column is a different seed.
Figure 7: Density percentage (%) of Att-Out, FF-1, and FF-2 masks for weight masking. Each row represents a different KG and each column is a different seed.
Figure 8: Venn diagrams for parameter overlap of three subnetworks identified under three different random seeds with weight masking, for each KG representation, location, and communication.
Figure 9: Venn diagrams for parameter overlap of three subnetworks identified under three different random seeds with input neuron masking, for each KG representation, location, and communication.
Figure 10: Jaccard similarity of different seed masks for the same KG with weight masking, (representation, location, and communication). The brighter the color, the higher the Intersection over Union.
Figure 11: Jaccard similarity of different seed masks for the same KG with input neuron masking, (representation, location, and communication). The brighter the color, the higher the Intersection over Union.
Figure 12: Venn diagrams for parameter overlap of three subnetworks identified under three different KGs with weight masking, for each seed 735, 1318, and 84.
Figure 13: Jaccard similarity of different KG masks for the same seed with weight masking, (735, 1318, and 84). The brighter the color, the higher the Intersection over Union.